# The Telephone Game: Evaluating Semantic Drift in Unified Models

### v1

@@ 3.1 MCD: Mean Cumulative Drift
where $S_{\delta}(g)$ is the average similarity at generation $g$ for distance mapping $\delta$, $M^{(g)}_{d,\delta}$ is the generated text or image at generation $g$, and $\operatorname{sim}$ denotes the similarity function from Tab. 1. To get overall drift, we compute mean across generations $S_{\delta}(g)$,
@@
3.2 SDR: Semantic Drift Rate
While MCD provides a holistic measure of semantic preservation, the Semantic Drift Rate (SDR) quantifies how rapidly a model’s outputs diverge from the original input. In MCD, a single distance mapping with high similarity values can dominate the overall score, potentially masking drift in other mappings. SDR addresses this by fitting a power curve across all directions, and obtain an averaged power curve, producing three parameters: $\alpha$, $\beta$, $\gamma$, which capture scaling, decay rate, and baseline similarity, respectively.
The power-law decay curve has the form $y=\alpha\cdot g^{-\beta}+\gamma$, where $y$ denotes similarity at generation $g$, $\alpha$ represents the extrapolated initial similarity, and $\gamma$ denotes the asymptotic baseline, the value at which a model eventually collapses to. Here, $\beta$ denotes the decay rate. $\beta$ is followed by a negative sign, which ensures that $g^{-b}$ decreases as $g$ increases, reflecting the fact that similarity diminishes over successive generations. More details on the power-law curve is provided in Appendix Fig. LABEL:fig:power_law. The power-law form is chosen because it models the empirically observed nonlinear decay: a sharp drop in early generations followed by a slower decline in later cycles.
To compute the SDR parameters, we use the dataset-level similarity scores $S_{\delta}(g)$ obtained in Eq. 1. We fit a power-law decay model, across the scores yielding the parameters $\alpha_{\delta}$, $\beta_{\delta}$, and $\gamma_{\delta}$ for each $\delta\in\mathcal{\Delta}$.
Finally we average each parameter across all distance mapping $\delta$ to obtain the final $\alpha$, $\beta$, and $\gamma$ which are used to construct the SDR power-law curve.
@@ 3.2 SDR: Semantic Drift Rate
3.3 MGG: Multi-Generation GenEval
To complement embedding-based similarities with object-level fidelity, we further extend GenEval [22] to our proposed multi-generation setting. The existing framework [22] is designed to assess text-to-image fidelity across multiple dimensions of quality. These dimensions include single_object, two_object, counting, colors, and positions, and attributes_binding.
For each task, GenEval proposes a diverse set of prompts such as ”a photo of a/an [COLOR] [OBJECT]”. Once a model has generated images for all the prompts, GenEval uses a pre-trained object detection model to detect and localize objects in the generated images. This process allows us to calculate the accuracy of the model for each task. An average of the task level accuracies is then denoted by GenEval overall accuracy. We build on the existing benchmark by incorporating the GenEval Rewritten dataset [6], adopting the newer OwlV2 object detection model [16], and extending evaluation across multiple generations. To calculate MGG, we first calculate the GenEval scores for each generation for all tasks. Then, similar to GenEval overall accuracy, we compute the tasks scores to obtain GenEval overall accuracy for each generation. Finally, we average the generation scores to obtain the MGG score. Higher MGG scores indicate better ability to produce semantically accurate and, context-preserving outputs.
@@ 4 Evaluations & Findings
(a) Text-First-Chain: $\text{text}^{(0)}\rightarrow\text{text}^{(g)}$ (MPNet)
(b) Text-First-Chain: $\text{text}^{(0)}\rightarrow\text{image}^{(g)}$ (CLIP)
(c) Image-First-Chain: $\text{image}^{(0)}\rightarrow\text{image}^{(g)}$ (DINO)
(d) Image-First-Chain: $\text{image}^{(0)}\rightarrow\text{text}^{(g)}$ (CLIP)
@@ 4.1 Datasets
For embedding based semantic drift analysis (MCD, SDR), we randomly sample 200 image-text pairs from each of the two challenging vision-language datasets, Nocaps [2] and DOCCI [17]. We denote this sample dataset as ND400. These corpora stress both novel objects and fine-grained details, making them well-suited to reveal drift that single-pass metrics do not capture. NoCaps introduces nearly 400 novel objects unseen in COCO and features more visually complex images. The novel objects enables testing models on out-of-domain. DOCCI was specifically curated to evaluate fine-grained reasoning in image-text models. The image captions cover attributes, spatial relationships, object counts, text rendering, and world knowledge. These data allow will allow us to evaluate models in their descriptive understanding or generation capabilities. For multi-generation GenEval evaluations (MGG), we employ the GenEval Rewritten dataset [6], which extends the short GenEval prompts into long descriptive texts which better match models’ outputs.
4.2 Evaluation of Unified Consistency
To conduct our evaluations, we construct three independent evaluation chains as shown in Tab. 1. Chain-1 (Exp. 1) is a Text-First-Chain that uses GenEval-Rewritten dataset. This chain is used exclusively for the multi‑generation GenEval score (MGG). Chain-2 (Exps. 2–4) is also a Text‑First-Chain constructed by initializing from the captions provided in ND400. Chain-3 (Exps. 5–7) follows the Image‑First-Chain and is initialized from the images in ND400. Chain-2 and Chain-3 are used to calculate the similarity $S$ used to compute our embedding‑based metrics: Mean Cumulative Drift (MCD) and Semantic Drift Rate (SDR).
4.3 Findings
@@
From our evaluations, we observe several interesting qualitative patterns. Fig. 5 illustrates six of such different ways in which unified models lose information under alternating $\texttt{T2I}\leftrightarrow\texttt{I2T}$ cycles: 1. Position Inconsistency: the model fails to preserve spatial relationships that are central to the scene, 2. Object Misidentification: low-fidelity renderings lead to incorrect re-captioning, 3. Style Transition: the model may change the style of an image, particularly for rare object pairings (e.g., a horse on a keyboard), 4. Quantity Inconsistency: numerical counts may be inflated, 5. Object Hallucinations: new elements are introduced, 6. Color Inconsistency: important colors are not retaine
Figure 7: Results for SDR obtained using the technique in Eq. 3.2 on both text-first and image-first settings. The results show a similar ranking across both settings. Although Show-o seems to be doing worse than Janus 1.3B in the text-first setup, the decay rate hints that it might perform better in the later generations.
Exp. Dataset Chain Direction Similarity Metric(s)
(start) ($\delta$) backbone
1 GenEval-R Text-First - - MGG
2 ND400 Text-First text $\rightarrow$ text MPNet MCD, SDR
3 ND400 Text-First text $\rightarrow$ text CLIP MCD, SDR
4 ND400 Text-First text $\rightarrow$ image CLIP MCD, SDR
5 ND400 Image-First image $\rightarrow$ image DINO MCD, SDR
6 ND400 Image-First image $\rightarrow$ image CLIP MCD, SDR
7 ND400 Image-First image $\rightarrow$ text CLIP MCD, SDR
Table 1: Summary of experimental settings. ND400 contains image-caption pairs, whereas GenEval Rewritten uses captions only. Each experiment is defined by its starting modality (text-first or image-first), the transformation direction ($\delta$), and the similarity backbone used for evaluation. CLIP for text-to-text, text-to-image, and image-to-image; DINO for image-to-image; and a MPNet for text-to-text. Experiment 1 corresponds to the multi-step GenEval configuration, and Experiments 2-7 capture semantic drift on ND400 through embedding-level similarity based metrics (MCD, SDR). GenEval-R: GenEval-Rewritten
Figure 8: MGG results on the GenEval Rewritten dataset. This heatmap shows the overall performance across the six tasks described in the GenEval [22] benchmark. On average, BAGEL consistently drifts the least from the semantic meaning of the original caption. In contrast, VILA-U and Janus 1.3B lose more than half of their first-generation score within just a few generations.
Figure 9: Comparison across MCD and MGG shows that BAGEL achieves the highest performance on both metrics, while VILA-U lags in both. LLaVA+SDXL and Janus 1.3B demonstrate strength in one metric but underperform in the other.
@@ 7 ND400 Image-First image $\rightarrow$ text CLIP MCD, SDR
Next, we present the empirical results in Fig. 6 which shows the scores obtained from Eq. 1 for all distance mappings, $\{\text{text}\!\to\!\text{text},\text{image}\!\to\!\text{text},\text{text}\!\to\!\text{image},\text{image}\!\to\!\text{image}\}$. These scores are later used to obtain both MCD and SDR. In the ideal case, the similarities should remain nearly constant across generations. Instead, as shown in these plots we observe consistent degradation in semantic fidelity, with modality dependent asymmetries. Plot 6(a) measure the similarity between the original caption and the text generated in Text-First-Chain. Top-performing models such as BAGEL and Blip-3o start with a high similarity ($\sim$ 0.70-0.75) and maintains it relatively well, ending around 0.60. In contrast, models like VILA-U and Janus 1.3B exhibit a much steeper decline, with VILA-U’s similarity dropping below 0.50, indicating that its generated texts or images quickly lose connection to the original prompt. Plot 6(b) and Plot 6(d) offers a cross-modal perspective, evaluating the $\text{text}\!\to\!\text{image}$, and $\text{image}\!\to\!\text{text}$ respectively. In both scenarios, BAGEL maintains a clear lead, while VILA-U’s generations drift so severely that their relevance to the original text becomes minimal at later stages. Across both plots, the overall model ranking at the last step is exactly same. Plot 6(c) measure visual fidelity by comparing the original image to the generated images at subsequent steps in Image-First-Chain. While the leading models perform similar to prior trends discussed above, we notice Janus 1.3B scoring high in the first generation (0.6), but eventually degrading to a low score in the last generation. Overall, this behavior of models performing well in the first generation, but eventually losing context along the generations is a characteristic not reliably captured by conventional single‑pass metrics.
Fig. 7, gives us a fine-grained look at the decay. The parameters used to plot these curves are shown in Tab. 3. The $\beta$ parameter tells us the decay rate for each model. BAGEL’s curves are visibly flatter, which is due to its lower $\beta$ value, indicating slower drift and greater cross‑modal stability. In contrast, LLaVA+SDXL and Janus 1.3B decline more steeply, reflecting faster semantic decay as shown in Fig. 7. Although Show-o appears weaker than Janus 1.3B in early text‑first steps, its slower decay rate ($\beta$) suggests relatively better behavior in later generations.
Fig. 8 reports MGG overall performance across categories. First generation scores are above 0.8 and >>0.95 for stronger models, which can mask qualitative differences; BAGEL, for instance, produces more faithful, higher‑quality objects than peers despite similar initial scores. As generations accumulate, the distribution widens and robustness differences emerge. As shown in Fig. 8, the results highlights that early quantitative parity does not guarantee long‑run stability: during Generations 1-2, Show-o attains overall GenEval and DINO similarities comparable to BAGEL, yet side‑by‑side inspection shows BAGEL is consistently more coherent and faithful. This divergence becomes numerically apparent only in later generations as Show-o’s decay accelerates, underscoring that cyclic evaluation reveals quality differences that single‑pass metrics can obscure. The results in separate GenEval tasks (Fig. 11) reveal a critical vulnerability in these models: their performance collapses most dramatically on compositional tasks like positioning and attribute binding. This may suggest weakness in composite attribute binding to be a cause for semantic drift.
Finally, Fig. 9 summarizes model performance by plotting MGG (y-axis) against $\texttt{MCD}_{avg}$ (x-axis), aligning embedding-level and object-level metrics. The rankings reveal a correlation between the two metrics for certain models (e.g., VILA-U, BAGEL). The decoupled LLaVA+SDXL system exhibits a unique performance profile, where it performs well across MGG, but struggles in MCD. It struggles to preserve the holistic semantics or the overall “vibe” of a scene, as reflected in its low embedding-level scores. However, its powerful generator allows it to perform comparatively better on object-level tasks, successfully rendering the primary subjects from a caption even as the broader context degrades. This hints at a disconnect between maintaining the content of an image and preserving its essential meaning. Across all evaluations, BAGEL consistently demonstrates the greatest resilience to semantic drift. This superior performance is likely due to its larger scale, as well as its architecture and training methodology. Further, BAGEL has a lot of carefully chosen architectural components, and it has been trained on a diverse set of filtered data, including various interleaved multimodal datasets. This may enable BAGEL to preserve important details, making it uniquely robust against compounding errors observed in our MGG evaluation framework.
@@ 5 Conclusion
We presented the Unified Consistency Framework (UCF), a cyclic evaluation that alternates image-to-text (I2T) and text-to-image (T2I) with complementary metrics (MCD for overall drift, SDR for decay rate, and MGG for multi-generation GenEval) to quantify how unified models preserve semantics over repeated modality shifts. Evaluating seven recent models on the sampled ND400 dataset shows substantial variability: BAGEL maintains the strongest cross-modal stability, VILA-U and JANUS variants drift quickly, and Show-o, while not always leading initially, degrades more gracefully across generations. These results demonstrate that single-pass benchmarks can overstate robustness, whereas cyclic evaluation reveals hidden inconsistencies between understanding and generation. We demonstrate that cyclic consistency is essential for reliable assessment, decay rates matter as much as initial scores.
Acknowledgements
@@
Appendix A Appendix
This appendix provides additional details and extended analyses that complement the results presented in the main paper. We first describe the models used in our experiments, including their parameterization and image generation settings. We then report further evaluations using CLIP embeddings, provide fitted parameters for our decay function, and present comprehensive results from the extended multi-generation GenEval benchmark.
A.1 Models & Parameters
Tab. 2 lists the models included in our evaluations, along with their parameter counts and image resolutions used during generation.
Name Parameters Image Resolution
BAGEL 14B - Mixture of Transformers $\times1024$
(7B Active)
@@ A.1 Models & Parameters
Table 2: Overview of models used in our experiments, including parameter counts and image resolution. The BAGEL model is a mixture-of-transformers architecture, where 7B parameters are active during inference.
A.2 Related Works

### v2

@@ 3.1 MCD: Mean Cumulative Drift
where $S_{\delta}(g)$ is the average similarity at generation $g$ for distance mapping $\delta$, $M^{(g)}_{d,\delta}$ is the generated text or image at generation $g$, and $\operatorname{sim}$ denotes the similarity function. To get overall drift, we compute mean across generations $S_{\delta}(g)$,
@@
3.2 MGG: Multi-Generation GenEval
To complement embedding-based similarities with object-level fidelity, we further extend GenEval Ghosh et al. (2023) to our proposed multi-generation setting. The existing protocol Ghosh et al. (2023) is designed to assess text-to-image fidelity across multiple dimensions of quality. These dimensions include single_object, two_object, counting, colors, and positions, and attributes_binding. For each task, GenEval proposes a diverse set of prompts such as ”a photo of a/an [COLOR] [OBJECT]”. Once a model has generated images for all the prompts, GenEval uses a pre-trained object detection model to detect and localize objects in the generated images. This process allows us to calculate the accuracy of the model for each task. An average of the task level accuracies is then denoted by GenEval overall accuracy. We build on the existing benchmark by incorporating the GenEval Rewritten dataset Chen et al. (2025), adopting the newer OwlV2 object detection model Minderer et al. (2024), and extending evaluation across multiple generations. To calculate MGG, we first calculate the GenEval scores for each generation for all tasks. Then, similar to GenEval overall accuracy, we compute the tasks scores to obtain GenEval overall accuracy for each generation. Finally, we average the generation scores to obtain the MGG score. Higher MGG scores indicate better ability to produce semantically accurate and, context-preserving outputs.
@@ 3.2 MGG: Multi-Generation GenEval
3.3 Single-Pass Human Evaluation (Cross-Consistency)
We complement our cyclic analysis with a single-step cross-consistency evaluation to highlight cross-modal fidelity issues. Given a ground-truth pair $(I,T)$, we first generate a caption $T^{(1)}=\mathrm{UM}_{\texttt{I2T}}(I)$ via I2T and an image $I^{(1)}=\mathrm{UM}_{\texttt{T2I}}(T)$ via T2I. We then assess whether $T^{(1)}$ and $I^{(1)}$ preserve the semantics of $(I,T)$ along two axes: (a) $I\rightarrow T^{(1)}$ consistency—does $T^{(1)}$ faithfully describe $I$? and (b) $T\rightarrow I^{(1)}$ consistency—does $I^{(1)}$ depict $T$? Multiple annotators independently evaluated each sample for fidelity and cross-consistency, and ranked model outputs by relevance to the original $(I,T)$ pair. For both directions, annotators rated outputs on a three-level scale: Good, Medium, or Poor. To ensure unbiased evaluation, generation-1 outputs from all models were displayed on a web interface with model identities hidden. Each sample page contained two sections: in the understanding section, annotators rated and ranked captions for the input image; in the generation section, they rated and ranked generated images for the input text prompt.
@@ 4 Evaluations & Findings
(a) Text-First-Chain: $\text{text}\rightarrow\text{text}$ (MPNet)
(b) Text-First-Chain: $\text{text}\rightarrow\text{image}$ (CLIP)
(c) Image-First-Chain: $\text{image}\rightarrow\text{image}$ (DINO)
(d) Image-First-Chain: $\text{image}\rightarrow\text{text}$ (CLIP)
@@
For embedding based semantic drift analysis (MCD), we randomly sample 200 image-text pairs from each of the two challenging vision-language datasets, Nocaps Agrawal et al. (2019) and DOCCI Onoe et al. (2024). We denote this sample dataset as Nocaps+Docci400. These corpora stress both novel objects and fine-grained details, making them well-suited to reveal drift that single-pass metrics do not capture. NoCaps introduces nearly 400 novel objects unseen in COCO and features more visually complex images. The novel objects enables testing models on out-of-domain. DOCCI was specifically curated to evaluate fine-grained reasoning in image-text models. The image captions cover attributes, spatial relationships, object counts, text rendering, and world knowledge. These data will allow us to evaluate models in their descriptive understanding or generation capabilities. For multi-generation GenEval evaluations (MGG), we employ the GenEval-R (GenEval Rewritten) dataset Chen et al. (2025), which extends the short GenEval prompts into long descriptive texts which better match models’ outputs.
4.1 Semantic Drift Protocol Findings
@@
From our evaluations, we observe several interesting qualitative patterns. Fig. 5 illustrates six of such different ways in which unified models lose information under alternating $\texttt{T2I}\leftrightarrow\texttt{I2T}$ cycles: 1. Position Inconsistency: the model fails to preserve spatial relationships that are central to the scene, 2. Object Misidentification: low-fidelity renderings lead to incorrect re-captioning, 3. Style Transition: the model may change the style of an image, particularly for rare object pairings (e.g., a horse on a keyboard), 4. Quantity Inconsistency: numerical counts may be inflated, 5. Object Hallucinations: new elements are introduced, 6. Color Inconsistency: important colors are not retained.
@@
Next, we present the empirical results in Fig. 6 which shows the scores obtained from Eq. 1 for all distance mappings, $\{\text{text}\!\to\!\text{text},\text{image}\!\to\!\text{text},\text{text}\!\to\!\text{image},\text{image}\!\to\!\text{image}\}$. These scores are later used to obtain MCD. In the ideal case, the similarities should remain nearly constant across generations. Instead, as shown in these plots we observe consistent degradation in semantic fidelity, with modality dependent asymmetries. Plot 6(a) measures the similarity between the original caption and the text generated in Text-First-Chain. Top performing models start with a high similarity ($\sim$ 0.65-0.70), however only BAGEL maintains it relatively well, ending around 0.65.
Figure 7: MGG results on the GenEval Rewritten dataset. This heatmap shows the overall performance across the six tasks described in the GenEval Ghosh et al. (2023) benchmark. On average, BAGEL consistently drifts the least from the semantic meaning of the original caption.
In contrast, models like VILA-U and Janus 1.3B exhibit a much steeper decline, with VILA-U’s similarity dropping below 0.40, indicating that its generated texts or images quickly lose connection to the original prompt. Plot 6(b) and Plot 6(d) offer a cross-modal perspective, evaluating the $\text{text}\!\to\!\text{image}$, and $\text{image}\!\to\!\text{text}$ respectively. In both scenarios, BAGEL maintains a clear lead, while VILA-U’s generations drift so severely that their relevance to the original text becomes minimal at later stages. Across both plots, the overall model ranking at the last step is exactly same. Plot 6(c) measures visual fidelity by comparing the original image to the generated images at subsequent steps in Image-First-Chain. While the leading models perform similar to prior trends discussed above, we notice Janus 1.3B scoring high in the first generation (0.6), but eventually degrading to a low score in the last generation. Overall, this behavior of models performing well in the first generation, but eventually losing context along the generations is a characteristic not reliably captured by conventional single‑pass metrics.
Figure 8: Comparison across MCD and MGG shows that BAGEL achieves the highest performance on both metrics, while VILA-U lags in both. The models align in a linear fashion, hinting at a correlation between the two scores.
Fig. 7 shows that while initial MGG scores are high, they can mask qualitative differences between models. For instance, BAGEL produces more faithful generations than Show-o even with similar initial scores, a divergence that only becomes numerically apparent in later generations as semantic drift occurs. This underscores that cyclic evaluation reveals quality differences that single-pass metrics obscure. Furthermore, performance collapses most dramatically on compositional tasks like positioning and attribute binding (Fig. 11), suggesting this weakness is a key cause of semantic drift. Overall performance, summarized in Fig. 8, plots MGG against $\texttt{MCD}_{avg}$ and reveals a correlation between object-level and embedding-level metrics. A notable exception is the decoupled LLaVA+SDXL system, which scores well on MGG but poorly on MCD, indicating it can render specific objects while failing to preserve holistic scene semantics. Across all evaluations, BAGEL consistently shows the most resilience to semantic drift, likely due to its scale, architecture, and training on diverse interleaved datasets, which makes it uniquely robust against the compounding errors our protocol exposes.
4.2 Human Evaluation Results
The results of human evaluations in Fig. 9 show that, unified models are comparatively better in image understanding than in generation. Due to this, most models show Poor-Generation-Good-Understanding type of cross-inconsistency. For each model, we also computed the mean ranking across all samples (lower values indicate better performance) to establish a human-based ordering. These rankings were then compared with our embedding-based metrics to evaluate the alignment between automated measures and human judgment. Fig. 12 illustrates the correlation between human rankings and the MCD metric. In the two subfigures, MCD scores are plotted against human rankings for generation (a) and understanding tasks. Both tasks yield a Pearson correlation of r<0.80r<-0.80 and an $R^{2}>0.65$, indicating a strong correlation and validating our embedding-based approach as a reliable proxy for human-perceived semantic consistency. In Fig. 13, we plot the GenEval score (a) and MGG score against human rankings in the image generation task. GenEval shows a Pearson correlation of r=0.753r=-0.753, while MGG exhibits an even stronger correlation with human rankings. This suggests that MGG, as a multi-generation metric, more accurately reflects model performance in line with human perception than the single-pass GenEval score.
Figure 9: Human evaluation of cross-consistency. First two plots show the percentage of samples (y-axis) rated with a fidelity score (color) for different models. We see that most models gain a high amount of Poor fidelity score in image generation, whereas understanding is pretty balanced, with Bagel almost always getting Medium or better. The third plot illustrates a finer look at the responses for the Bagel model. We see that while Bagel has 10% of Good-understanding-Bad-Generation type of inconsistency, it does not have any of the other (Bad-Understanding-Good-Generation) kind of inconsistency.
@@ 5 Conclusion
We introduced the Semantic Drift Protocol (SDP), a cyclic evaluation framework that alternates image-to-text (I2T) and text-to-image (T2I) to measure how unified models preserve meaning over repeated modality shifts. By combining embedding-based metrics (MCD) and object-level fidelity (MGG), SDP exposes vulnerabilities that single-pass evaluations cannot capture. Evaluating seven recent models on the sampled Nocaps+Docci400 dataset shows substantial variability: BAGEL maintains the strongest cross-modal stability, VILA-U and JANUS variants drift quickly, and Show-o, while not always leading initially, degrades more gracefully across generations. Human evaluations confirm these findings, showing that automated metrics like MCD strongly align with human judgments. These results demonstrate that single-pass benchmarks can overstate robustness, whereas our cyclic evaluation validated by human judgment reveals hidden inconsistencies between image understanding and image generation. We conclude that cyclic evaluation is essential for reliable assessment of unified models.
6 Acknowledgment
@@
Appendix
This appendix provides additional details and extended analyses that complement the results presented in the main paper. We first describe the models used in our experiments, including their parameterization and image generation settings. We then report further evaluations using CLIP embeddings, and present comprehensive results from the extended multi-generation GenEval benchmark.
Appendix A Models & Parameters
Name Parameters Res.
BAGEL 14B – Mixture of Transformers (7B Active) $\times1024$
@@ Appendix A Models & Parameters
Tab. 1 lists the models included in our evaluations, along with their parameter counts and image resolutions used during generation. The BAGEL model is a mixture-of-transformers architecture, where 7B parameters are active during inference.
Appendix B Related Works
Unified Models T2I generation has advanced with diffusion-based models such as DALL·E 2 Ramesh et al. (2022), Imagen Saharia et al. (2022), and Stable Diffusion Rombach et al. (2022), which synthesize high-fidelity images from textual prompts. Image captioning, on the other hand, has evolved from CNN-RNN pipelines Shi et al. (2015) to transformer-based decoders Cornia et al. (2020); Liu et al. (2023) trained with large web-scale data. Recent works in unified models have started investigating how to unite understanding and generation under one architecture. Chameleon Team (2025) is one of the early works in this domain which aimed to auto-regressively generate text tokens and image embeddings. Later, Transfusion Zhou et al. (2024) fused the auto-regressive and diffusion loss within a single architecture. Show-o Xie et al. (2024) has also used two different objectives, next token prediction for text generation, and masked token prediction Chang et al. (2022) for image generation. VILA-U Wu et al. (2025) uses next token prediction with different text and vision decoders. Janus and Janus-pro Wu et al. (2024) employ separate encoders for image input during understanding and generation. The idea is that a model might require different level of information for understanding and generation. Other works like BLIP3-o Chen et al. (2025) demonstrates good quality of image generation by leveraging a separate diffusion transformer head. A recent work, BAGEL Deng et al. (2025) demonstrates some unique capabilities of unified models by training on a large-scale interleaved dataset.